# Test-negative designs with various reasons for testing: statistical bias and solution

### v2

@@ Discussion
As TNDs become increasingly popular for evaluating VE based on observational data, the demand for valid, robust, and appropriate statistical analyses for TNDs is also growing. In this paper, we focus on an emerging issue that a TND may recruit participants with varying reasons for testing, and clarify the VE estimands associated with the various reasons. Of course, the suggested approach relies on collecting accurate information on the reason for testing at the time of recruitment participation. We note that severe bias can be caused if the reasons for testing are ignored. In summary, TNDs with various reasons for testing bring not only challenges in analyses but also present opportunities in obtaining precise estimates of VE on different populations. Not discussed here is the possibility of using estimates of $VE_{s}(\boldsymbol{x})$ and $VE(\boldsymbol{x})$ to estimate of $RR_{P}$, the relative risk for progression to symptoms associated with vaccination. Our statistical analysis and proposed methods provide a handle to take advantage of this opportunity and can potentially inform future TND studies.
@@
In designing and implementing TND studies, it is thus important to understand why people are tested. If this information is missing, an estimated VE can be biased. In this case, the magnitude and direction of this bias is likely intractable absent further information. Therefore, when tests for asymptomatic subjects are expected, we recommend recording the reasons for testing in patient enrollment or in the surveillance system. For this purpose, an example questionnaire used by the California Department of Public Health in the United States is available in the Supplementary Material of [7].
@@
Our statistical framework relies on the assumption that a subject should have a single, accurate reason for testing. In practice, a patient may have two or more reasons for testing or, on the contrary, unclear or misclassified reasons for testing. While this may not be an issue for mandatory testing such as case contact tracing, bias may occur with self-reporting patients. This issue can be mitigated by adopting survey methods to promote accuracy and honesty [24] and can also be examined by a sensitivity analysis. We leave this topic for future research.
@@
To derive the VE measures given various reasons for testing, we have mainly focused on a simple setting with a binary vaccination status and a binary outcome. In the Covid-19 setting, however, the VE may differ by vaccination doses, the definition of outcomes (infected, diseased, hospitalized, or dead), and the variant of infection. With multi-level $V$ and $I$, our framework can be applied to evaluate the effect of each specific vaccination on each type of outcomes separately, provided that vaccination and outcomes are well recorded and the assumptions in Table 1 hold. On this topic, [25] proposes measuring symptom severity to alleviate selection bias on the odds ratio scale; non-binary healthcare-seeking behavior is handled as well. A comprehensive examination of these different scenarios is an important future direction.

### v3

@@ Discussion
As TNDs become increasingly popular for evaluating VE based on observational data, the demand for valid, robust, and appropriate statistical analyses for TNDs is also growing. In this paper, we focus on an emerging issue that a TND may recruit participants with varying reasons for testing, and clarify the VE estimands associated with the various reasons. Of course, the suggested approach relies on collecting accurate information on the reason for testing at the time of recruitment participation. We note that severe bias can be caused if the reasons for testing are ignored. In summary, TNDs with various reasons for testing bring not only challenges in analyses but also present opportunities for obtaining precise estimates of VE on different populations. Our statistical analysis and proposed methods provide a handle to take advantage of this opportunity and can potentially inform future TND studies.
@@
In designing and implementing TND studies, it is thus important to understand why people are tested. If this information is missing, an estimated VE can be biased. In this case, the magnitude and direction of this bias are likely intractable absent further information. Therefore, when tests for asymptomatic subjects are expected, we recommend recording the reasons for testing in patient enrollment or in the surveillance system. For this purpose, an example questionnaire used by the California Department of Public Health in the United States is available in the Supplementary Material of [7].
@@
Our statistical framework assumes that a subject should have a single, accurate reason for testing. In practice, however, a patient may have two or more reasons for testing or, on the contrary, unclear or misclassified reasons for testing. While this may not be an issue for mandatory testing such as case contact tracing, bias may occur with self-reporting patients. This issue can be mitigated by adopting survey methods to promote accuracy and honesty [22] and can also be examined by a sensitivity analysis. We leave this topic for future research.
In order to remove the collider bias on testing, an important assumption we implicitly assume is that the observed covariates include all necessary variables for achieving the conditional independence between $I$ and $(R=r,T=1)$. Similar to the no-unmeasured confounder assumption in causal inference, evaluating or relaxing this assumption requires additional data or framework. Recently, [23] proposed sensitivity analysis for TNDs, which can be adapted to examine how the results are impacted without this assumption.
@@
To derive the VE measures given various reasons for testing, we have mainly focused on a simple setting with a binary vaccination status, a binary outcome, and a binary healthcare-seeking behavior. In the Covid-19 setting, however, the VE may differ by vaccination doses, the definition of outcomes (infected, diseased, hospitalized, or dead), and the variant of infection. With multi-level $V$ and $I$, our framework can be applied to evaluate the effect of each specific vaccination on each type of outcomes separately, provided that vaccination and outcomes are well recorded and the assumptions in Table 1 hold. Furthermore, the healthcare-seeking behavior should reflect the patients’ propensity to seek care, which is unlikely to be captured by a binary variable. Therefore, the selection bias from testing may be reduced instead of fully removed by conditioning on a binary healthcare-seeking behavior in classical TNDs [24, 5]. On this topic, [25] proposes measuring symptom severity to alleviate selection bias on the odds ratio scale; non-binary healthcare-seeking behavior is handled as well. In addition, [26] used negative controls to address the bias from healthcare-seeking behavior. A comprehensive examination of these different scenarios is an important future direction.